What is the pH of a sodium formate solution prepared by adding 0.680 grams of sodium formate to 100.0 ml of water at 25.0°C? The Ka at 25.0°C for formic acid is 1.8 × 10-4.



1
Expert's answer
2021-11-22T10:18:04-0500

Solution:

sodium formate = HCOONa


Calculate the moles of sodium formate:

The molar mass of HCOONa is 68 g/mol

Hence,

(0.680 g HCOONa) × (1 mol HCOONa / 68 g HCOONa) = 0.01 mol HCOONa


Calculate the molar concentration (CM) of a sodium formate solution:

CM(HCOONa) = moles of sodium formate / volume of solution

CM(HCOONa) = (0.01 mol / 0.1 L) = 0.1 mol/L = 0.1 M

CM(HCOO) = CM(HCOONa) = 0.1 M


Сalculate the dissociation constant:

K = Kw / Ka

K = (1.0×10-14) / (1.8×10-4) = 5.56×10-11


The equilibrium reaction for dissociation of is,

HCOONa(aq) + H2O(l) ⇌ HCOOH(aq) + NaOH(aq)

HCOO(aq) + H2O(l) ⇌ HCOOH(aq) + OH(aq)


ICE Table:



[HCOO] = 0.1 - x

[HCOOH] = [OH] = x


K = [OH] × [HCOOH] / [HCOO]

K = (x]) × (x) / (0.1 - x) = (x2) / (0.1 - x) = 5.56×10-11

By solving the terms, we get:

x = 2.358×10-6

[OH] = x = 2.358×10-6 M


From [OH], we can calculate pOH:

pOH = -log[OH] = -log(2.358×10-6) = 5.627

 

From pOH, we can calculate pH:

pH + pOH = 14;

pH = 14 - pOH = 14 - 5.627 = 8.373 = 8.37

pH = 8.37

 

Answer: pH = 8.37